Calculate the cost of debt from the following information: company cost of capital = 15%, cost of equity capital = 16 per cent, D = $0.2 million, and V = $0.5 million.


A. 12.5%
B. 13.5%
C. 10.5%
D. 15.5%
Answer #1
Solution:
Answer is B. 13.5%
Working Notes:
cost of capital is WACC =15%
Cost of equity Ke = 16%
Cost of debt Kd = ???
Value of Debt D = $0.2 million
Value of Company V = $0.5 million
Value of Equity E = V -D = $0.5 million - $ 0.2 million
Value of Equity E = V -D = $0.3 million
WACC = (E/V)Ke + (D/V) Kd
15% = (0.3/0.5) x 16% + (0.2/0.5) x Kd
Kd= (15%- 9.6%) x 0.5/0.2
Kd= 5.4% x 0.5/0.2
Kd= 13.5%
cost of debt Kd = 13.5%
Notes: WACC is the weighted average cost of capital which is also called cost of capital.
Please feel free to ask if anything about above solution in comment section of the question.